A 6-foot tall man is standing near a tree on level ground(the shadow of the tree is 18 feet long). If the man's shadow is 4 feet long, how many feet tall is the tree?
Answer by AnlytcPhil(1806)

A 6-foot tall man is standing near a tree on level ground(the shadow of the
tree is 18 feet long).  If the man's shadow is 4 feet long, how many feet tall
is the tree? 

---------------------------------------------

Man's height IS TO Man's shadow AS Tree's height IS TO Tree's shadow

Let the tree's height be x. Then we have

                   6 IS TO 4 AS x IS TO 18

Replace the "IS TO"'s by " / " and the "AS" by " = "

                     6/4 = x/18

Cross multiply:

                    4x = 6(18)
                    4x = 108
                     x = 27 feet

Edwin